# A Decade of Human-Robot Interaction through Immersive Lenses: A Literature Review on Extended Reality as a Research Instrument in Social Robotics

### v1

@@ 4.4 RQ4: Characteristics of Researchers and Participants
From the researcher perspective, social XR-HRI is dominated by technology-oriented disciplines, which account for roughly 70% of author backgrounds. Interdisciplinary, psychological, and other non-technical fields (e.g., education, medicine, design, law) appear only sporadically. The field is strongly shaped by early-career scholars: PhD candidates (n = 58) and students (n = 34) make up the majority, supported by substantial senior involvement from professors (n = 57). Only three contributors are software engineers. The binary sex distribution is uneven, with 92 men and 70 women. Since sex was inferred rather than self-reported, it refers to observable sex rather than gender identity. Geographically, the field is concentrated in Western countries—led by the USA (n = 13) and Germany (n = 8) – with smaller contributions from several European countries. Asian participation is scattered, and the Global South is nearly absent, leaving culture-specific interaction norms underrepresented. A similar pattern emerges on the side of participants (overall age: M = 31.16, SD = 20.82). The sample is predominantly male (490 vs. 335 female, 3 diverse), with 255 unspecified cases. Most participants are university students (616), making recruitment convenient but unrepresentative for real-world user groups. Children (n = 53), teenagers (n = 38), and older adults (n = 42) appear only when tied to application domains such as education or healthcare. Age distribution ranges widely – from primary school age to over 75 – but median values fall within the student bracket. Study locations mirror this bias: most sessions took place in university laboratories (n = 653), compared to far fewer in public spaces (n = 76) or school settings (n = 73). Documentation gaps remain substantial: nearly one-third of participant information (266 entries) lacks details on identity, origin, or age.
@@ 5 Discussion
The present literature review shows that XR technologies are currently used primarily as controllable laboratory instruments in HRI research, with a focus on evaluation and simulation. While this focus enables precise experimental control and cost-effective iteration, it comes at the expense of ecological validity and long-term practical transfer. Domains such as public space, industry, and hospitality are currently severely underrepresented, even though these contexts are likely to be crucial for the integration of social robots. Technical and methodological bottlenecks characterize the status quo. On a technical level, limitations in head-mounted displays, tracking accuracy, network stability, and a lack of haptic feedback hinder realism. However, these obstacles are expected to be resolved with the advent of new XR devices. Methodologically, short lab sessions, small homogeneous samples, and self-reported measurements dominate. Objective sensor technology and automated behavioral analysis remain the exception. These limitations mean that social XR-HRI results are often only partially generalizable. Researchers need to critically reflect on why they use XR for research and at the same time remember what individual benefits the medium brings with it. Another key shortcoming is the lack of diversity – both in the research teams and among the participants. The majority of studies are conducted in Western-influenced, technology-oriented contexts with predominantly young adults. Children, older adults, and people from non-Western cultures are rarely found in the samples. Interdisciplinary approaches that systematically integrate social, psychological, or ethical perspectives are rare. In terms of content, most studies operate at an interaction level that is stimulus-driven rather than dialogical. Language is frequently used, but almost exclusively in script-based form. Multimodal, adaptively responsive robot interactions remain a research shortcoming. Particularly striking is the lack of important reporting, e.g., on the hardware and software used, but also on robots. These are essential for working in accordance with scientific ethics, but also for conducting reproducible and clean research. A driving factor is that, to date, there is no uniform XR taxonomy for describing such things. The use of XR is described in many ways, but rarely classified with precision. Existing taxonomy approaches (e.g.,  [1]) offer valuable insights, but have not yet been tailored to the specifics of HRI. The need for a domain-specific social XR-HRI taxonomy that clearly structures roles, contexts, and forms of use is particularly noteworthy, thus offering researchers a clear classification option.
@@ 6 Future Roadmap for social XR-HRI
We propose a five-part roadmap for XR as a research instrument for social HRI in order to establish the medium as a reliable research and transfer instrument for social robots:
@@
    Strengthen Application Contexts – Many social XR-HRI studies are still conducted in digital laboratory settings rather than in specific application contexts. Yet social HRI is shaped by context and may not translate well from simplified lab reproductions [63]. Researchers should take advantage of the greater ecological validity of XR and conduct their studies in realistic application environments, using their own creations or ready-made scenes (e.g., from the Unity Store). This approach also promotes a certain degree of replicability, as virtual environments can be shared with the community.
@@
    Build More Robust and Testable Technological Iterations – The SLR shows that current social XR-HRI systems often remain proof-of-concept prototypes, with WoZ control and rudimentary interactions of the robot (e.g., scripted dialogue). While this reflects the very high implementation effort for XR applications and, at the same time, there is pressure regarding the structural characteristics of scientific research to have very short publication cycles, the field would benefit from longer-term and iterative systems that go beyond “flat” demos. Sustainable foundations must be laid, e.g., through toolkits, modular XR–robot integration frameworks, and authoring tools with reusable interaction modules. This process could reduce overhead and improve robustness and reproducibility.
@@
    Foster Methodological Innovation – The focus should shift from predominantly subjective self-reports toward multimodal, objective data (e.g., automated logs, real-time biosignals). Although modern XR headsets technically support measures such as eye tracking, integrated physiological sensing remains rare in social XR-HRI. Standardized, easy-to-use solutions (e.g., Unity/Unreal plugins) could lower barriers and enable reliable incorporation of physiological metrics into study designs.
@@
    Embed Diversity in Samples and Research Teams – Social XR-HRI studies often rely on WEIRD, technology-oriented participant pools and similarly homogeneous research teams, which limits generalizability and carries the risk of systematic bias [61]. These circumstances naturally pose a challenge for the entire research landscape, not just social XR-HRI. Nevertheless, researchers should expand recruitment across cultures, age groups, genders, and technical experience levels, reflecting diversity in team composition and participatory design practices. The key question is how to achieve this diverse sample. We suggest that recruitment via platforms such as Prolific – especially for online surveys – or the targeted engagement of communities via associations and networks could be one way forward. Furthermore, awareness of diversity issues is a cornerstone that needs to be strengthened among researchers in order to increase their willingness to include a broad spectrum of people in their own studies.
  5. 5.
    Build Reporting Standards – A key problem in XR-HRI research is the lack of standardized, reporting standards that capture both technical details and HRI-relevant aspects. While conceptual taxonomies exist, they mainly categorize systems and interaction paradigms rather than defining what must be reported for transparency and reproducibility [65]. A report-oriented taxonomy that translates such conceptual frameworks into a structured documentation scheme for social XR-HRI is missing. The SLR shows that crucial information is often incomplete or omitted. This lack of transparency undermines reproducibility and interpretability, making it difficult to judge whether effects stem from interaction design, robot behavior, measurement choices, or implementation differences. Because XR setups and robot platforms are highly heterogeneous, a multidimensional reporting standard is needed, covering XR roles, technical factors, contextual factors, and methodological characteristics. A practical solution would be a checklist-based reporting framework (PRISMA-style) tailored to XR-HRI. Standardized reporting would improve cross-study comparability, enable replication, and make study designs, robot behaviors, and data pipelines more transparent and reusable.
@@ 7 Limitations
The SLR has several limitations. First, it should be noted that the search strategy only covered four databases, even if these are major publication venues for HRI. This could mean that relevant articles were overlooked. Second, the coding process involves subjective interpretations, especially in the inductive categories. Third, the review focuses on studies published since 2015; while this corresponds to the availability of modern XR systems, it ignores earlier and most recent work. Fourth, this SLR is limited in its significance because the studies examined lack information, particularly on the demographic characteristics of the participants, the hardware/software specifications, and the methodological details in some categories.
@@ 8 Conclusion
This systematic literature review provides an initial overview of how XR has been used as a research tool in social robotics over the past ten years. The results show that the unique possibilities of XR are not being fully exploited. Rigid and unimodal robot interactions often dominate, with robots frequently presented as objects to be observed in laboratory studies, limited use of AR and MR, low participant diversity, and a strong reliance on self-reported measurements. While these approaches offer a high degree of experimental control, they come at the expense of ecological validity, inclusivity, and transferability to the real world. In particular, technical and methodological bottlenecks have been identified, as well as inconsistent reporting practices that prevent reproducibility and comparability. To take social XR-HRI a step further towards being a recognized research tool, future work should focus on long-term studies, place more trust in the functional capabilities of virtual robots, promote diversity among researchers, but above all among participants, and create a standardized social XR-HRI taxonomy. Doing so will not only enhance the scientific robustness of social XR-HRI research but also ensure that its insights and applications are relevant, transferable, and socially inclusive.
@@ 8.0.1 Acknowledgements
The presented work was supported by the RuhrBots competence center (16SV8693) funded by the Federal Ministry of Research, Technology and Space Germany. We would like to thank Alexandra-Michelle Drechsler for providing feedback on the manuscript.

### v2

@@ 4.4 RQ4: Characteristics of Researchers and Participants
A total of 162 researchers were identified as authors. The binary sex distribution is uneven, with 92 men and 70 women. Since sex was inferred rather than self-reported, it refers to observable sex rather than gender identity. In terms of the background of researchers involved, social XR-HRI is dominated by technology-oriented disciplines, which account for roughly 70% of author backgrounds (n=81). Interdisciplinary, psychological, and other non-technical fields (e.g., education, medicine, design, law) appear only sporadically (n=43). The field is strongly shaped by early-career scholars: PhD candidates (n = 58) and students (n = 34) make up the majority, supported by substantial senior involvement from professors (n = 57). Only three contributors are software engineers. Geographically, the field is concentrated in Western countries – led by the USA (n = 13) and Germany (n = 8) – with smaller contributions from several European countries. Asian participation is scattered, and the Global South is nearly absent, leaving culture-specific interaction norms underrepresented.
A similar pattern emerges on the side of participants (overall age: M = 31.16, SD = 20.82). The sample is predominantly male (490 vs. 335 female, 3 diverse), with 255 unspecified cases. Most participants are university students (n = 616), making recruitment convenient but unrepresentative for real-world user groups. Children (n = 53), teenagers (n = 38), and older adults (n = 42) appear only when tied to application domains such as education or healthcare. Age distribution ranges widely – from primary school age to over 75 – but median values fall within the student bracket. Study locations mirror this bias: most sessions took place in university laboratories (n = 653), compared to far fewer in public spaces (n = 76) or school settings (n = 73). Documentation gaps remain substantial: nearly one-third of participant information reports (n = 266) lacks details on identity, sex/originx educational status, origin, or age.
@@ 5 Discussion
The present literature review shows that XR technologies are currently used primarily as controllable laboratory instruments in HRI research, with a focus on evaluation and simulation. While this focus enables precise experimental control and cost-effective iteration, it comes at the expense of ecological validity and long-term practical transfer. Domains such as public space, industry, and medical facilities are currently severely underrepresented, even though these contexts are likely to be crucial for the integration of social robots. Technical and methodological bottlenecks characterize the status quo. On a technical level, limitations in head-mounted displays, tracking accuracy, network stability, and a lack of haptic feedback hinder realism. However, these obstacles are expected to be resolved with the advent of new XR devices. Methodologically, short lab sessions, small homogeneous samples, and self-reported measurements dominate. Objective sensor technology and automated behavioral analysis remain the exception. These limitations indicate that social XR-HRI results are often only partially generalizable. Researchers should critically reflect on their reasons for using XR in research while considering the specific benefits the medium provides. Another key shortcoming is the lack of diversity – both in the research teams and among the participants. Recent work has concluded that limited sample diversity can lead to blind spots in HRI research [61, 15]. On the researcher team side, increased diversity can lead to improved performance [31] and reduce bias against marginalized groups [51]. The majority of studies are conducted in Western-influenced, technology-oriented contexts with predominantly young adults. Children, older adults, and people from non-Western cultures are rarely found in the samples. Interdisciplinary approaches that systematically integrate social, psychological, or ethical perspectives are likely rare, presumably due to their limited number. In terms of content, most studies operate at an interaction level that is stimulus-driven rather than dialogical. Language is frequently used, but almost exclusively in script-based form. Multimodal, adaptively responsive robot interactions remain a research shortcoming. Particularly striking is the lack of important reporting, e.g., on the hardware and software used, but also on robots. These are essential for working in accordance with scientific ethics, but also for conducting reproducible research. A driving factor is that, to date, there is no uniform XR taxonomy for describing such specifications. The use of XR is described in many ways, but rarely classified with precision. Existing taxonomy approaches (e.g., [1]) offer valuable insights, but have not yet been tailored to the specifics of HRI. The need for a domain-specific, social XR-HRI taxonomy that clearly structures roles, contexts, and forms of use is particularly noteworthy, as it would offer researchers a clear classification option/framework.
@@ 6 Future Roadmap for Social XR-HRI
We propose a four-part roadmap for XR as a research instrument for social HRI in order to establish the medium as a reliable research and transfer instrument for social robots:
@@
    Strengthen Application Contexts. Many social XR-HRI studies are still conducted in digital laboratory settings rather than in specific application contexts. This means that robots and their interactions are often viewed in isolation. Yet, social HRI is shaped by context and may not translate well from simplified lab reproductions [24]. Researchers should take advantage of the greater ecological validity of XR and conduct their studies in realistic application environments, using their own creations or ready-made scenes (e.g., from the Unity Store). This approach also promotes a certain degree of replicability, as virtual environments can be shared with the community.
@@
    Build More Robust and Testable Technological Iterations. The present review shows that current social XR-HRI systems often remain proof-of-concept prototypes, with WoZ control and rudimentary interactions of the robot (e.g., scripted dialogue). While this reflects the very high implementation effort required for XR applications and the pressure of short publication cycles in scientific research, the field would benefit from longer-term, iterative systems that go beyond “flat” demos. Sustainable foundations must be laid, e.g., through toolkits, modular XR–robot integration frameworks, and authoring tools with reusable interaction modules. This process could reduce overhead and improve robustness and reproducibility.
@@
    Embed Diversity in Samples and Research Teams. Social XR-HRI studies often rely on WEIRD, technology-oriented participant samples and similarly homogeneous research teams, which limits generalizability and carries the risk of systematic bias [61]. These circumstances naturally pose a challenge for the entire research landscape, not just social XR-HRI. Nevertheless, researchers should expand recruitment across cultures, age groups, genders, and technical experience levels, reflecting diversity in team composition and participatory design practices. We argue that a key lever is stronger collaboration across labs, disciplines, and sectors – especially through international collaborations that can provide local contextual expertise, improve cultural validity, and enable access to participant groups that are difficult to reach. Furthermore, awareness of diversity issues is a cornerstone that needs to be strengthened among researchers in order to increase their willingness to include a broad spectrum of people in their own studies.
@@
    Build Reporting Standards. A key problem in XR-HRI research is the lack of reporting standards that capture both technical details and HRI-relevant aspects. While conceptual taxonomies exist, they mainly categorize systems and interaction paradigms rather than defining what must be reported for transparency and reproducibility [63]. The SLR shows that crucial information is often incomplete or omitted. This lack of transparency undermines reproducibility and interpretability, making it difficult to judge whether effects stem from interaction design, robot behavior, measurement choices, or implementation differences. Because XR setups and robot platforms are highly heterogeneous, a multidimensional reporting standard is needed, covering XR roles, technical factors, contextual factors, and methodological characteristics. A practical solution could be a checklist-based reporting framework (PRISMA-style) tailored to XR-HRI. Standardized reporting would improve cross-study comparability, enable replication, and make study designs, robot behaviors, and data pipelines more transparent and reusable.
@@ 7 Limitations
The SLR has several limitations. First, it should be noted that the search strategy only covered four databases, even if these are major publication venues for HRI and have a certain technical focus. Second, the coding process involves subjective interpretations, especially in the few inductive categories. However, qualitative coding aims to derive concepts and categories through interpretation rather than purely objective classification. Third, the review focuses on studies published since 2015; while this corresponds to the availability of modern XR systems, it ignores earlier and most recent work. Fourth, this review is limited in its precision of details because the examined studies lack information, particularly on the demographic characteristics of the participants, the hardware/software specifications, and the methodological details in some categories. Furthermore, the order of authors listed in publications does not reveal who contributed how much to the individual works. This could distort the evaluation in terms of the distribution of disciplines. Fifth, literature reviews are often subject to publication bias, meaning that most studies with significant results are published, while others may be omitted. This could be exacerbated in the field of XR because the medium is relatively new, highly experimental, and complex. The review presented may therefore not cover all studies that would normally have been relevant.
@@ 8 Conclusion
This literature review provides an overview of how XR has been used as a research tool in social robotics over the past ten years. The results show that the unique possibilities of XR are not being fully exploited. Rigid and unimodal robot interactions often dominate, with robots frequently presented as objects to be observed in laboratory studies, limited use of AR and MR, low participant diversity, and a strong reliance on self-reported measurements. While these approaches offer a high degree of experimental control, they come at the expense of ecological validity, inclusivity, and transferability to the real world. In particular, technical and methodological bottlenecks have been identified, as well as inconsistent reporting practices that prevent reproducibility and comparability. To take social XR-HRI a step further towards being a recognized research tool, future work should focus on long-term studies, place more trust in the functional capabilities of virtual robots, promote diversity among researchers – but above all among participants – and create a standardized social XR-HRI taxonomy. Doing so will not only enhance the scientific robustness of social XR-HRI research but also ensure that its insights and applications are relevant, transferable, and socially inclusive.
@@ 8.0.1 Acknowledgements
The presented work was supported by the RuhrBots competence center (16SV8693) funded by the German Federal Ministry of Research, Technology and Space. We would like to thank Alexandra-Michelle Drechsler, Lukas Erle and Ramón-Darius Imort for providing feedback on the manuscript.